NEED HELP SOLVING:
BERMUDA TRIANGE HAS A PERIMETER OF 3075MI. THE SHORTEST SIDE MEASURES 75MI LESS THAN THE MIDDLE SIDE, AND THE LONGEST SIDE MEASURES 375MI MORE THAN THE MIDDLE SIDE. FIND THE LENGTHS OF THE THREE SIDES.

Answer by Grilly-O(3)

1) Let x= the middle side; x-75= the short side; x +375= the long side
2) For the equation, since you are dealing with the perimeter, we will just add the sides together to equal the given perimeter. So the equation will be:
x + (x-75) + (x +375)= 3075
3) Next, you just solve that equation for x. So, x=925. That is the middle lenght. The Short side is 925-75=850. The long side will be 925 +375= 1300.
4) The lengths of the sides are 850 miles, 925 miles, and 1,300 miles.